For point charge 10 µC and point charge -4 µC located at the same positions as in the previous question (5 m and 4 m, respectively), determine the magnitude of the net electric field E at the origin (in N/C).

Your answer should be a number with two decimal places, do not include the unit.

Hint: 1µC = 10-6C
Answer #1

The first point charge g1 10 C 10* 10-6 C is located at T1=5 m

Electric field at origin due to charge q_1 is Ч1 E1 4тєот , negative sign indicates that it is directed towards left. (away from charge q_1)

The second point charge 42 =-4 C -4 10-C is located at r2=4 m

Electric field at origin due to charge q_2  is 42 E2 4TED , positive sign indicates that it is directed towards right. (towards the charge q_2)

Net electric field at origin due to the two charges is  91 42 E-+Ε- .2 4περπ 4περα3

10 10 4 10 4(8.85 10-12)(5)? 4(8.85 10-12) (4)2 6 E =

E 1348.77 N/C 1.35 103 N/C

Magnitude of electric field at the origin is E 1348.77 N/C 1.35 103 N/C